Assume that a bond will make payments every six months as shown on the following timeline: 21 Cash flow $25.00 $25.00 $25.00 $1,025.00 a. What is the maturity of the bond (in years)? b. What is the coupon rate (in percent)? c. What is the face value? a. What is the maturity of the bond (in years)? The maturity of the bond in years is years. (Round to the nearest integer.) b. What is the coupon rate (in percent)? The coupon rate is 96 (Round to two decimal places.) c. What is the face value? The face value is (Round to the nearest dollar)
Answer #1

(a) Maturity of the bond in years = ( 6 months) (total payments) = (6/12)(22) = (132/12) = 11 years . Because payment is made every six months. It implies that in one year two payments are made.

(b) Coupon payment =( coupon rate)(face value) / Number of coupons per year)

Coupon payment = 20

Face value = 1025

Number of coupons per year = 2

Coupon rate = (22)(2) /1025 = 44/1025 = 0.0429 = 4.29%

(c) Face value = $1025